OpenStudy (anonymous):

solve equation and check solution -3(t-6)-(t+8)=9

OpenStudy (nincompoop):

perform the distributive property

OpenStudy (anonymous):

use distribute property that is like multiply -3(t-6) and then subtract remain t + 8 then subtract the number not with no variable then divide to get t by it's self

OpenStudy (anonymous):

ok I am still lost

OpenStudy (anonymous):

distribute what is -3(t-6) (Hint multiply both by -3)

OpenStudy (anonymous):

what's your answer.... for that part?

OpenStudy (anonymous):

-3t+18

OpenStudy (anonymous):

ok since it's subtraction by - sybol -3t +18 -t - 8 combine terms

OpenStudy (anonymous):

you get -4t +10= 9 // -10 on both sides and divide to get t

OpenStudy (anonymous):

ok I am new to this and I am 56 years old dude,lol. I will say 4t+10=9 is 4t+10-10=9-10 4t=-1 t=-1/4 maybe

OpenStudy (anonymous):

yes !!!!!!!!

OpenStudy (anonymous):

SWEET!!!! You are awesome thank you so much, have a great night.